Hi, can I please have your help with the following worded equation?
James is twice as old as Andy. Twelves years ago, James was three times as old. How old are they now?
I just need help getting it into an equation
Thank you
Sally

Answer by mananth(16946)

Andy age = x years
James = 2x
..
12 years ago their ages were
andy x-12
James 2x-12
..
2x-12= 3(x-12)
2x-12 = 3x -36
add 12
2x-12+12 = 3x-36+12
2x= 3x-24
add -3x
2x-3x= -3x+3x-24
-x = -24
divide by 1
x=24
Andy's age
James age = 48